# Multi-Objective Neural Network Assisted Design Optimization of Soft Fin-Ray Grippers for Enhanced Grasping Performance

### v1

@@ III MLP Model
A multilayer perceptron (MLP) is chosen to handle complex nonlinear behavior between the input and output features using intermediate hidden layers. In this feedforward neural network, the mean squared error is considered as the loss for the fitted model. The input layer of the MLP has three neurons consisting of three features: the thickness of the front and supporting beams, the thickness of the crossing beams, and the equal spacing between each crossbeam. The output layer consists of four layers, including maximum contact forces and maximum displacements of the tip in the x- and y-directions. The MLP has three hidden and dropout layers. We use 80% of the data for training, 10% for the validation and the remaining 10% for the test. The number of neurons in each hidden layer, the type of activation function for hidden layers, the regularization rate, and the dropout rate are hyperparameters that are searched through the specified range as summarized in Table II. In this search, a specific parameter value has higher scores when it performs better in the validation set. Moreover, the output layer’s activation function is set as Sigmoid function. The final MLP architecture and the selected hyperparameters are summarized in Table III. Subsequently, K-Fold Cross-Validation is used to ensure that our model is trained and tested on representative samples, reducing bias and enhancing the overall performance.
TABLE II: Grid search over specified parameter values
@@
Hidden layers activation function types Leaky ReLU-Sigmoid-Tanh
Regularization rates 0-0.001-0.01
Dropout rates 0-0.1-0.2
TABLE III: Final MLP architecture
@@
Number of neurons in the third hidden layer 7
Hidden layers activation function type Tanh
Regularization rate 0
Dropout rate 0
@@
where $N$ is the number of data points, $\hat{y}$ is the predicted value of $y$, $y_{i}$ is the values of $y$ variable in the sample, and $\bar{y}$ is the mean value of $y$. The mean squared error (MSE), mean absolute error (MAE) and $R^{2}$ score for training, validation, and test data of the target features are tabulated in Table IV. We note that features $F_{x}$ and $F_{y}$, appearing on the table, represent the contact forces in the x- and y-directions, respectively. In a similar vein, features $D_{x}$ and $D_{y}$ indicate the tip displacements, respectively, in the x- and y- directions. These numbers suggest that the model is suitably generalized to unseen data and our model has a balance between bias and variance, avoiding overfitting and underfitting. Therefore, based on the performance metrics and numerical results presented here, the MLP is capable of precise predicting of the target features for various internal structures.
TABLE IV: MSE, MAE, and $R^{2}$ score performance metrics
Error Type Feature MSE MAE $R^{2}$ score
Training $F_{x}$ 0.00084 0.02245 0.99
Validation $F_{x}$ 0.00068 0.01954 0.99
Test $F_{x}$ 0.00104 0.02537 0.99
Training $F_{y}$ 0.00351 0.04156 0.94
Validation $F_{y}$ 0.00217 0.03760 0.94
Test $F_{y}$ 0.00143 0.02804 0.98
Training $D_{x}$ 0.00338 0.04282 0.94
Validation $D_{x}$ 0.00177 0.03141 0.95
Test $D_{x}$ 0.00254 0.04111 0.95
Training $D_{y}$ 0.00241 0.03562 0.96
Validation $D_{y}$ 0.00184 0.02823 0.96
Test $D_{y}$ 0.00198 0.03707 0.97
IV Multi-objective Design Optimization
We employ a multi-objective optimization algorithm to find the optimal internal structure for the Fin-Ray finger. We use NSGA-II multi-objective genetic algorithm, which generates offspring by using a particular kind of crossover and mutation method. The selection process for the next generations is based on comparisons of the non-dominated sorting and crowding distances[21]. Three design parameters are considered in the dimensional optimization, including the thickness of the front and supporting beams, the thickness of the crossing beams, and the equal spacing between the crossbeams, which are considered input features for the MLP in the previous section. These design parameters are constrained in the range that the MLP is trained according to Table I. The MLP model developed in the previous section is used here for accurate prediction of our objectives.
Two objectives of optimization are readily determined from the four target features of the MLP using Eqs. 5 and 6:
@@ IV Multi-objective Design Optimization
where $F$ and $D$ are the predicted magnitude of the maximum contact force and magnitude of the maximum tip displacement, $F_{x}$ and $D_{x}$ are the predicted maximum force and maximum tip displacement in the x-direction, and $F_{y}$ and $D_{y}$ are the predicted maximum force and maximum tip displacement in the y-direction, respectively. As aforementioned, these magnitudes are considered as our two objectives. The maximum tip displacement towards the grasped object is regarded a scale for how suitable a design is at handling delicate objects and highlights the Fin-Ray finger’s adaptability to the cylindrical objects. A higher value for the contact force resembles the tendency of the finger to possess a higher load capacity. Achieving a large grasping force demands enhanced rigidity, which is in contrast with the first objective and it is against the softness, necessary for delicate grasping.
After the determination of the objectives and design parameters, NSGA-II is utilized to find the optimal design of the Fin-Ray finger. The Pymoo library allows to implement the NSGA-II algorithm by defining our problem with the specific parameters mentioned in Table V[22]. The set of optimal solutions, which is also known as the Pareto front, is presented in Fig. 5. Pareto front narrows down the solutions to a set in which we can make choice in a trade-off of one objective over the others. Depending on application, one of these solutions is selected to achieve desirable design.
TABLE V: NSGA-II algorithm parameters
@@
We also conduct a series of numerical experiments, in which we feed the MLP random numbers and compare the predicted target features with the Pareto front. However, the Pareto front is not dominated by the experiment data or dataset as illustrated in Fig. 6.
Figure 6: Experiment data, dataset, and Pareto front
V Conclusion
@@
In this research, we have presented an comprehensive approach for designing the Fin-Ray finger to enhance the grasping performance. A finite element framework for predicting the forces during the gripper’s interaction with cylindrical objects has been proposed. Using finite element simulations, we have developed a dataset consisting of distinct Fin-Ray internal structures. A model has been forged through training the neural network. The chosen MLP hyperparameters have been selected via a grid search over the specified range of values. This MLP model has been used to find the optimal structures with the NSGA-II algorithm. The set of all Pareto efficient solutions has been found and validated with the experiment data and the dataset itself. Hence, it is found that the present methodology achieved the desired results for designing the optimal Fin-ray finger, assisting to make a choice in the trade-off of one objective over the other.

### v2

@@ 3 MLP Model
where $r$ is the correlation coefficient, $x_{i}$ are the values of the variable $x$ in the sample, and $\bar{x_{i}}$ is the mean of $x$. Also, $y_{i}$ denotes values of the variable $y$ in the sample, and $\bar{y_{i}}$ is the mean of $y$ [4]. This heatmap approach suggests that the correlations between each input and target feature are significant, which shows the suitability of the selected features. Other exploratory data analysis (EDA) techniques are used to reduce the effect of any outlier and address similar problems. Hence, we organize the data to ensure accuracy, consistency, and suitability for the task in a proper manner.
A multilayer perceptron is chosen to handle complex nonlinear behavior between the input and output features using intermediate hidden layers. In this feedforward neural network, the mean squared error is considered as the loss for the fitted model. The input layer of the MLP has three neurons consisting of three features: the thickness of the front and support beams, the thickness of the crossbeams, and the equal spacing between each crossbeam. The output layer consists of four neurons corresponding to the maximum contact forces and tip displacements in the x and y-directions, which are $F_{x}$, $F_{y}$, $D_{x}$, and $D_{y}$. The MLP has three hidden and dropout layers. We use 80% of the data for training, 10% for validation and the remaining 10% for testing. Using this search technique, we tried to find the best combination of hyperparameters. Hence, the number of neurons in each hidden layer and the type of activation function for hidden layers are hyperparameters that are searched over the specified values as summarized in Table 2.
Table 2: Grid search over specified parameter values
@@
Hidden layers activation function types ReLU-Sigmoid-Tanh
After the grid search, we trained the final MLP architecture with the Adam optimizer using a learning rate of 0.001, batch size of 1, and 50 epochs. Furthermore, K-Fold cross-validation is used to tackle possible overfitting problems. This approach yields lower validation errors and smoother Pareto search.
Figure 5: Training and validation loss curves
In this search, a specific parameter value has higher scores when it performs better in the validation set. Moreover, the output layer’s activation function is set to the sigmoid function to hard-encode the physics that all design variables are bounded and non-negative.. The final MLP architecture and the selected hyperparameters are summarized in Table 3. Subsequently, K-Fold cross-validation is used to ensure that our model is trained and tested on representative samples, reducing bias and enhancing the overall performance.
Table 3: Final MLP architecture
@@
Number of neurons in the third hidden layer 9
Hidden layers activation function type ReLU
@@
where $N$ is the number of data points, $\hat{y}$ indicates the predicted value of $y$, $y_{i}$ is the values of $y$ variable in the sample, and $\bar{y}$ denotes the mean value of $y$. The mean squared error (MSE), mean absolute error (MAE) and $R^{2}$ score for test data of the target features are tabulated in Table 4. We note that features $F_{x}$ and $F_{y}$, appearing on the table, represent the contact forces in the x- and y-directions, respectively. In a similar vein, features $D_{x}$ and $D_{y}$ indicate the tip displacements, respectively, in the x- and y- directions. These numbers suggest that the model is suitably generalized to unseen data and our model has a balance between bias and variance, avoiding overfitting. Therefore, based on the performance metrics and numerical results presented here, the MLP is capable of predicting of the target features for various internal structures.
Table 4: MSE, MAE, and $R^{2}$ score performance metrics for test set
Feature MSE MAE $R^{2}$ score
Fx 0.00055 0.01823 0.99
Fy 0.00350 0.04569 0.96
Dx 0.00320 0.04958 0.94
Dy 0.00152 0.03419 0.97
4 Multi-objective Design Optimization
We employ a multi-objective optimization algorithm to find the internal structure for the Fin-Ray finger. We use NSGA-II, which generates offspring by using a particular kind of crossover and mutation method. The selection process for the next generations is based on comparisons of the non-dominated sorting and crowding distances [7]. Three design parameters are considered in the dimensional optimization, including the thickness of the front and support beams, the thickness of the crossbeams, and the equal spacing between the crossbeams, which are considered input features for the MLP in the previous section. These design parameters are constrained to the range in which the MLP is trained, as given in Table 1. The MLP model developed in the previous section is used here for accurate prediction of our objectives. Two optimization objectives are readily determined from the four target features of the MLP:
@@ 4 Multi-objective Design Optimization
where $F$ and $D$ represent the predicted magnitudes of the maximum contact force and magnitude of the maximum tip displacement. Also, $F_{x}$ and $D_{x}$ are, respectively, the predicted maximum force and maximum tip displacement in the x-direction, and $F_{y}$ and $D_{y}$ are the corresponding quantities in the y-direction. As aforementioned, these magnitudes are considered as our two objectives. The maximum tip displacement towards the grasped object is regarded a measure for how suitable a design is at handling delicate objects and highlights the Fin-Ray finger’s adaptability to the cylindrical objects. A higher value for the contact force indicates the tendency of the finger to possess a higher load capacity. Achieving a large grasping force demands enhanced rigidity, which is in contrast with the first objective and it is opposed to the softness necessary for delicate grasping.
After the determination of the objectives and design parameters, NSGA-II is utilized to find the design of the Fin-Ray finger. The Pymoo library allows us to implement the NSGA-II algorithm by defining our problem with the specific parameters mentioned in Table 5 [3]. We choose population size of 500 to maintain diversity, and we give sufficient evolution depth without excessive cost using 100 generations. The optimization termination criterion is considered as a maximum of 50 generations. No specific optimization constraint is defined for the optimization problem. The three input variables are normalized between 0 and 1, and their real bounds are limited to the geometric specifications. A crossover rate of 0.9 pushes mixing, while a mutation rate of 0.1 brings diversity without over-perturbing designs. These numbers are aligned with common practice and lead to satisfactory results here. For other optimization variables, parameters, and methods, the Pymoo default setup is used.
Table 5: NSGA-II algorithm parameters
@@
The set of optimal solutions, which is also known as the Pareto front, is presented in Fig. 6. The Pareto front narrows down the solutions to a set in which we can make a choice in the trade-off between objectives. Depending on the application, one of these solutions is selected to achieve a desirable design.
@@
We also generate a series of random points, feed them into the MLP, and compare the predicted target features with the Pareto front. As illustrated in Fig. 7, the Pareto front is not dominated by these synthetic samples. With the limited amount of training data available, the MLP showed a slight tendency to underestimate the Fin-Ray forces and displacements compared to the dataset points, particularly near the regions of the Pareto front. However, it mainly reproduced nearby trade-offs. This effect can be attributed to the regression toward the mean, which can be observed in neural network models trained with limited data. Consequently, the predicted forces and displacements along the Pareto front are generally lower than those obtained from FEM simulations for the same geometries.
Figure 7: Random MLP Samples, dataset, and Pareto front
In order to examine the Pareto front results, we select three points from it and compare them with the FEM simulation results. These three points and their corresponding Fin-Ray finger geometric parameters are presented in Table 6. Depending on the intended application of the Fin-Ray gripper, the chosen design may vary. Here, Points A and B represent the designs with the highest tip displacement and the highest force, respectively. In addition to these points, Point C is considered a balanced solution. Since the displacement and force values along the Pareto front have different ranges, these values were normalized between 0 and 1. Among the normalized Pareto front points, the one with the minimum Euclidean distance to the (1,1) is selected as Point C. The FEM simulation results, Pareto front data, and corresponding errors are also summarized in Table 7. It is expected to achieve higher contact forces when the overall stiffness of the Fin-Ray structure increases. Relatively, this can be addressed when the thickness of different parts of the Fin-Ray gets larger. In this case, the resulting Fin-Ray would not be suitable for delicate tasks. However, according to Table 6, it is shown that font and support beam thicknesses have larger impact on the objectives, while crossbeam thickness is less impactful. Regarding tip displacement, this comparison demonstrates a relatively low error between the FEM and Pareto front outputs. The error increases for the force column. This behavior is somehow expected because forces are more sensitive to local stiffness changes and nonlinear effects than tip displacements and small differences in the predicted deformation can lead to more noticeable differences in the reaction forces. On the other hand, given the limited FEM dataset, the predictive model naturally prioritizes the smoother and dominant trends in the input and output mapping. Tip displacements vary relatively smoothly with the geometric parameters, whereas contact forces exhibit stronger nonlinearities, especially in the high-force region as depicted in Fig. 4. Therefore, the predictive model makes lower relative errors for displacements compared to forces. Nevertheless, these errors remain within an acceptable range for the design parameters and trade-off analysis targeting in this work.
Table 6: Geometry properties of selected Pareto front structures
Point Front and Support Crossbeams Spacing Distance
Beams Thicknesses (mm) Thickness (mm) ($m$)
A 1.500 1.600 10.000
B 4.000 1.600 11.320
C 3.310 1.599 10.000
Table 7: Comparison of Pareto front and FEM simulations results
Point Pareto Front Results FEM Results Error (%)
Displacement Force Displacement Force Displacement Force
A 31.609 mm 17.065 N 33.223 mm 16.032 N 4.857 6.443
B 21.668 mm 86.536 N 21.837 mm 94.143 N 0.777 8.080
C 27.845 mm 66.583 N 28.196 mm 61.276 N 1.246 8.661
5 Conclusion
@@
In this research, we have presented a comprehensive approach for designing Fin-Ray fingers to enhance grasping performance. A finite element framework for predicting the forces during the finger’s interaction with cylindrical objects has been proposed. Using finite element simulations, we have developed a dataset consisting of 120 simulation results of the tip displacements and the contact forces for distinct Fin-ray internal structures. A model has been forged through the training of the neural network. The MLP hyperparameters have been selected by a grid search over the specified range of values. This MLP model has been used to find the designs with the NSGA-II algorithm. The set of all Pareto efficient solutions has been found and validated with the random data generated by the MLP. Despite differences in the predicted force and displacement values by the MLP model compared to the simulation output of the same geometry, the optimization framework identified geometries with respect to the defined objectives. This shows that the surrogate model is sufficiently accurate to produce valid designs and show trade-offs. To assess the results of the optimization process, three different points from the Pareto front have been selected, and using their geometric properties, we have simulated the corresponding Fin-Ray finger structures. According to the simulation results of these fingers, we have compared the results to the Pareto front data. This comparison shows a promising relation between FEM simulations and design optimization solutions. However, performing experiments on selected Fin-Ray finger structures is suggested as future studies to validate the results. In addition, in this study we have considered the magnitude of tip displacement and contact force as the design objectives to enhance grasping performance. Moreover, our simulations have been done with only a certain geometry for the grasped object. In addition, future work should focus on dataset improvements to increase surrogate reliability. As a future study, our methodology can be generalized to more complicated geometries, better representatives for design objectives, different object sizes and materials as well. In conclusion, it is found that the present methodology achieved the desired results for designing the Fin-ray finger, assisting in making a choice between objectives.